# A Lagrangian Approach to the Inhomogeneous Incompressible Euler Equation

### v2

@@ 4 Lagrangian Analyticity of Solution
$\mathsf{Tr}(\nabla\ell^{(m)})=\nabla\cdot\ell^{(m)}=\sum_{k=2}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m,r_{i}\geq 1}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})}):=W^{(m)}$ (64)
Here the expansion eventually leads to a polynomial of lower order $\nabla\ell^{(r)}$s’ enries with degree at most $m$.
@@
$\nabla\cdot\ell^{(m+1)}=\frac{1}{m+1}\nabla\cdot(\rho_{0}^{-1}(\nabla p^{(m)}-U^{(m)}))=W^{(m+1)}$
@@
$\nabla p^{(m)}=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}+\nabla\cdot(\rho_{0}^{-1}U^{(m)})]$ (65)
@@
We aim to show that the formal series (66) converges absolutely.
Here, for later convenience, we decompose $U^{(m)}=\mathbf{P}U^{(m)}+\nabla\varphi^{(m)}$. Notice then
@@
First, we need to show that $\ell^{(m+1)}\in H^{s+1}$, given $\{\ell^{(k)}\}_{k=1}^{m}\subset H^{s+1}$. Note that with our assumption (A), $\Delta_{\rho_{0}}$ is uniformly elliptic. The above equality (65) reads
@@
$\ell^{(m+1)}=\frac{1}{(m+1)\rho_{0}}\big[-\mathbf{P}U^{(m)}+\nabla(p^{(m)}-\varphi^{(m)})\big]$ (68)
@@
Therefore, $\mathbf{P}U^{(m)}\in H^{s+1}$ will follow if we succeed in showing for any $1\leq k\leq m$:
$\mathbf{P}[\nabla^{*}\ell^{(k)}(\ell^{(m+1-k)}\rho_{0})]\in H^{s+1},\quad\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]\in H^{s+1}$
Following from combining Lemma 4.2 and a simple manipulation on the second term:
$\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]=-\mathbf{P}[\nabla(\ell^{(m+1-k)}\cdot\ell^{(k)})\rho_{0}]=-\mathbf{P}[\nabla^{*}\ell^{(m+1-k)}\cdot\ell^{(k)}\rho_{0}]-\mathbf{P}[\nabla^{*}\ell^{(k)}\cdot\ell^{(m+1-k)}\rho_{0}]$
@@
$\lVert\mathbf{P}U^{(m)}\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}$ (69)
@@
$\nabla(p^{(m)}-\varphi^{(m)})=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})]$
@@
$\lVert\nabla(p^{(m)}-\varphi^{(m)})\rVert_{H^{s+1}}\leq C(\rho_{0})(m+1)\lVert W^{(m+1)}\rVert_{H^{s}}+\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}$
Since $W^{(m+1)}$ consists of the sum of product of $\nabla\ell^{(r)}$ with $1\leq r\leq m$, algebra property of $H^{s}$ implies $W^{(m+1)}\in H^{s}$, with:
$\lVert W^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}$ (70)
@@
$\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}$
Hence we conclude from (68):
$\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}+\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}$ (71)
@@
$\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\leq\sum_{j=2}^{d}C_{j}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}$ (72)
@@
$\lvert\Psi_{t}^{\prime}(\zeta)\rvert\leq\sum_{j=2}^{d}jC_{j}\zeta^{d-1}=1-J^{\prime}(\zeta)\leq 1-\delta,\quad\text{for all }\zeta\in[0,\alpha_{*}]$
Hence for each $t$, $\Psi_{t}(\zeta)$ is a contraction. Moreover, we notice that $\Psi_{t}$ is increasing, hence:
@@
The implicit function theorem implies that locally around $(0,0)$, there exists an analytic function $\zeta_{*}(t)$ such that $P(t,\zeta_{*}(t))=0$. By uniqueness of solution of $P(t,\cdot)$, we conclude the formal series solution must coincide with analytic function $\zeta_{*}(t)$, which concludes the absolute convergence of series $\tilde{\zeta}$ in $t\in(0,T(\alpha^{*}))$, hence also the convergence of $\bar{\zeta}(t)$. Hence, Proof of the Lagrangian analyticity of IIE with convergence radius $t\in(0,T(\alpha^{*}))$ is complete. ∎

### v3

@@ 4 Lagrangian Analyticity of Inhomogeneous Incompressible Fluids
    $\mathsf{Tr}(\nabla\ell^{(m)})=\nabla\cdot\ell^{(m)}=\sum_{k=2}^{m}\frac{(-1)^{k}}{k}\sum_{r_{1}+\cdots+r_{k}=m,r_{i}\geq 1}\mathsf{Tr}(\nabla\ell^{(r_{1})}\cdots\nabla\ell^{(r_{k})})=:W^{(m)}$ (61)
    Here, the expansion eventually leads to a polynomial of lower order $\nabla\ell^{(r)}$s’ entries with degree at most $m$.
@@
    $\nabla\cdot\ell^{(m+1)}=\frac{1}{m+1}\nabla\cdot(\rho_{0}^{-1}(\nabla p^{(m)}-U^{(m)}))=W^{(m+1)},$
@@
    $\nabla p^{(m)}=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}+\nabla\cdot(\rho_{0}^{-1}U^{(m)})].$ (62)
@@
    We aim to show that the series (63) converges absolutely.
    Here, we consider the Leray-Hodge decomposition $U^{(m)}=\mathbf{P}U^{(m)}+\nabla\varphi^{(m)}$. Notice then
@@
    First, we need to show that $\ell^{(m+1)}\in H^{s+1}$, given $\{\ell^{(k)}\}_{k=1}^{m}\subset H^{s+1}$. Note that with our assumption 2.1, $\Delta_{\rho_{0}}$ is uniformly elliptic. The above equality (62) reads
@@
    $\ell^{(m+1)}=\frac{1}{(m+1)\rho_{0}}\big[-\mathbf{P}U^{(m)}+\nabla(p^{(m)}-\varphi^{(m)})\big].$ (65)
@@
    Therefore, $\mathbf{P}U^{(m)}\in H^{s+1}$ will hold true if we succeed in showing for any $1\leq k\leq m$:
    $\mathbf{P}[\nabla^{*}\ell^{(k)}(\ell^{(m+1-k)}\rho_{0})]\in H^{s+1},\quad\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]\in H^{s+1}.$
    Combining Lemma 4.2 and a simple manipulation on the second term:
    $\mathbf{P}[\nabla\rho_{0}(\ell^{(m+1-k)}\cdot\ell^{(k)})]=-\mathbf{P}[\nabla(\ell^{(m+1-k)}\cdot\ell^{(k)})\rho_{0}]$
    $\mathbf{P}[\nabla^{*}\ell^{(m+1-k)}\cdot\ell^{(k)}\rho_{0}]-\mathbf{P}[\nabla^{*}\ell^{(k)}\cdot\ell^{(m+1-k)}\rho_{0}],$
@@
    $\lVert\mathbf{P}U^{(m)}\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}.$ (66)
@@
    $\nabla(p^{(m)}-\varphi^{(m)})=\nabla\Delta_{\rho_{0}}^{-1}[(m+1)W^{(m+1)}-\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})].$
@@
    $\lVert\nabla(p^{(m)}-\varphi^{(m)})\rVert_{H^{s+1}}\leq C(\rho_{0})(m+1)\lVert W^{(m+1)}\rVert_{H^{s}}+\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}.$
    Since $W^{(m+1)}$ consists of the sum of product of $\nabla\ell^{(r)}$ with $1\leq r\leq m$, the fact that $H^{s}$ is an algebra implies $W^{(m+1)}\in H^{s}$, with:
    $\lVert W^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1,r_{i}\geq 1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}},$ (67)
@@
    $\lVert\nabla\Delta_{\rho_{0}}^{-1}\nabla\cdot(\rho_{0}^{-1}\mathbf{P}U^{(m)})\rVert_{H^{s+1}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}.$
    Hence, we conclude from (65):
    $\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\lesssim\sum_{k+j=m+1;k,j\geq 1}\lVert\nabla\ell^{(k)}\rVert_{H^{s}}\lVert\nabla\ell^{(j)}\rVert_{H^{s}}+\sum_{j=2}^{d}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}},$ (68)
@@
    $\lVert\nabla\ell^{(m+1)}\rVert_{H^{s}}\leq\sum_{j=2}^{d}C_{j}\sum_{r_{1}+\cdots+r_{j}=m+1}\prod_{i=1}^{j}\lVert\nabla\ell^{(r_{i})}\rVert_{H^{s}}.$ (69)
@@
    $\lvert\Psi_{t}^{\prime}(\zeta)\rvert\leq\sum_{j=2}^{d}jC_{j}\zeta^{j-1}=1-J^{\prime}(\zeta)\leq 1-\delta,\quad\text{for all }\zeta\in[0,\alpha_{*}]$
    Hence, for each $t$, $\Psi_{t}(\zeta)$ is a contraction. Moreover, we notice that $\Psi_{t}$ is increasing, therefore:
@@
    The implicit function theorem implies that locally around $(0,0)$, there exists an analytic function $\zeta_{*}(t)$ such that $P(t,\zeta_{*}(t))=0$. By uniqueness of root of $P(t,\cdot)$, we conclude the formal series solution must coincide with analytic function $\zeta_{*}(t)$, which concludes the absolute convergence of series $\tilde{\zeta}$ in $t\in(0,T(\alpha^{*}))$, hence also the convergence of $\bar{\zeta}(t)$. Hence, proof of the Lagrangian analyticity of IIE with convergence radius $t\in(0,T(\alpha^{*}))$ is complete.